Shift report 20011125 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Nov 25 2001 - 07:25:53 EST

  • Next message: Apache: "Shift report 20011125 08:00-16:00"

    ------------------------------------------------------------
    Date: Sun Nov 25 07:25:53 2001, Shift: 00:00-08:00
    Supervisor: Andrei Makeev, with: 
    Log book#11 pages 32 - 34
    
    Shift Summary:
    ------------------------------------------------------------
    1. After Tandem foil has been replaced, beam
    came back, with useful hours from 4 a.m. till
    6 a.m. From that time RHIC is preparing for 22 GeV
    running.
    
    2. 0ne run (#5983) with 110k trig.6 events.
    When run #5983 ended, I turned down all the
    voltages in the DCs as FV asked. The TPCs are
    in "parking mode".
    Runs #5984, 5985 are DVM runs for Ramiro.
    
    3. Magnets control program gives incorrect readuot
    of -3000.09 A for the D5 current. The actual
    current setting in D5 is correct (was checked
    by MCR person).
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Nov 25 2001 - 07:26:36 EST